OpenStudy (blackops2_143):

Evaluate. 2 – 4 • 3 + 6 • 3 A. 8 B. 12 C. 0 D. –28

563blackghost (563blackghost):

Use PEMDAS.... So we would first do multiplication...

OpenStudy (123ab456c):

do you want the answer or do u want me to explain it

OpenStudy (blackops2_143):

can you explain it?

563blackghost (563blackghost):

So first \[4 \times 3 = ?\]and \[6 \times 3=?\]

OpenStudy (blackops2_143):

12 and 18

563blackghost (563blackghost):

Correct ^^ So we have this now.... \[2-12+18\]Now we would subtract 2 and 12 since we go from left right.... so... \[2-12=?\]

OpenStudy (123ab456c):

well your answer is 8 i will show my work 2−(4)(3)+(6)(3)

OpenStudy (blackops2_143):

10?

OpenStudy (blackops2_143):

thx

OpenStudy (123ab456c):

wait are you saying thanks to me

563blackghost (563blackghost):

Actually its -10 ^^ So we now add -10 with 18....\[-10 + 18 = ?\]This will get you your answer...

OpenStudy (blackops2_143):

yes both of you

OpenStudy (123ab456c):

no its not i thought that to!:) i did this a fw years back

OpenStudy (123ab456c):

can you fan me and medal me i would really appreciate that!:)

563blackghost (563blackghost):

Hope I was of help ^^ thanx for medal...If you ever need help in the future just tag me in your question....

OpenStudy (123ab456c):

2 – 4 • 3 + 6 • 3 i will show my work 2−(4)(3)+(6)(3) =2−12+(6)(3) =−10+(6)(3) =−10+18 =8

OpenStudy (123ab456c):

do u get how its not -10

563blackghost (563blackghost):

Hmmm.....it is negative 10 otherwise we would not get the total of 8...

OpenStudy (123ab456c):

EXACLY now ur learning!!!! ;0;0

563blackghost (563blackghost):

but you stated "do you get how its "NOT" -10"....